OpenStudy (anonymous):

Find the horizontal asymptote (if any) of the graph of the following function. y= -6x^4-x+6/3x^6-7x+6 please help

OpenStudy (anonymous):

Try to think about what happens if you put in a large value for \(x\), say, \(x=10^6\). On top, you'd get \(-6*10^{24}-10^6+6 \)which is really close to \(-6*10^{24}\). On the bottom, we similarly get something really close to \(3*10^{36}\). Then the whole fraction is very close to \[\frac{-6*10^{24}}{3*10^{36}}=\frac{-2}{1000000000000}.\] I hope the point I'm trying to make is clear.

OpenStudy (anonymous):

i dont really understand couldnt the 6's cancel?

OpenStudy (anonymous):

They actually don't because we can only cancel multiplicative factors in fractions. Let me see if I can find a link to the method I was attempting to illustrate.

OpenStudy (anonymous):

Ah, here's a nice summary of the properties of rational functions: https://people.richland.edu/james/lecture/m116/polynomials/rational.html

OpenStudy (anonymous):

alright ty man im pretty sure i can do it now